OpenStudy (anonymous):

whats the integral of 1/ ( x * (3-x) ) I know it should be a ln ( something)

OpenStudy (lalaly):

do u know hw to use partial fractions?

OpenStudy (lalaly):

\[\frac{1}{x(3-x)} = \frac{A}{x}+\frac{B}{3-x}\]

OpenStudy (lalaly):

\[1=A(3-x)+bx\]\[1=3A-Ax+Bx\]\[3A=1\]\[A=\frac{1}{3}\]\[-A+B=0\]\[\frac{-1}{3}+B=0\]so\[B=\frac{1}{3}\]

OpenStudy (lalaly):

\[\frac{1}{x(3-x)}=\frac{1}{3x} + \frac{1}{3(3-x)}\]

OpenStudy (anonymous):

Thanks i can integrate that.

OpenStudy (lalaly):

:)